# Compensation behaviour in trilayered anisotropic 6-state clock model

### v1

@@ 3.2 Anisotropic regime: $D>0$
In this section, we explore the anisotropic response of the clock spins. By varying the anisotropy parameter D, we systematically investigate the evolution of key observables such as magnetisation, susceptibility and specific heat. Figure-7(a) shows the thermal variation of total magnetisation as function of temperature for different anisotropy values. As temperature decreases, magnetisation grows, exhibiting a clear phase transition at $T_{c}$. Below $T_{c}$, the total magnetisation again vanishes at a lower temperature, denoted as the compensation point $T_{comp}$, indicating the compensation behaviour. The critical temperature $T_{c}$ and compensation temperature $T_{comp}$ are precisely obtained from the peak positions of magnetic susceptibility curves Figure-7(b). The corresponding specific heat profiles further corroborate these findings by exhibiting peaks near both temperatures.
@@
Now, if we increase D, the susceptibility peak shifts toward the right, indicating the critical temperature increases. $T_{comp}$ also shifts slightly. This is captured in the phase diagram in D-T plane shown in Figure-9. This diagram summarizes the evolution of both the thermodynamic phase transition temperature $(T_{c})$ and the compesation temperature $(T_{comp})$ as a function of anisotropy $D$. A monotonic increase in $T_{c}$ observed with increasing $D$, indicationg that stronger single-site anisotropy stabilizes the ordered phase and shifts the thermodynamic transition to higher temperatures. In contrast, $T_{comp}$ exhibit only a slight upword trend with increasing $D$ and remains well separated from $T_{c}$ across the probed range. The critical temperature $T_{c}$ is prescribed by collective critical fluctuations, which are highly sensitive to the anisotropy parameter $D$. Increasing $D$ enhances spin alignment and supressess thermal disorder, thereby stabilizing the ordered phase and shifting $T_{c}$ to higher temperatures. Besides, the compensation temeperature $T_{comp}$ does not correspond to bonafide thermodynamic singularity. Instead it arises from a subtle cancellation between the magnetizations of individual sublattices. This balance dependes primarily on competing intra and inter layer exchange interactions and local spin configurations rather than on long-range critical fluctuations, resulting in only a weak variation of $T_{comp}$ with D.
@@
We also analyze the sublattice magnetization of different layers for different strengths of anisotropy. As positive $D$ influences magnetic ordering along $x$ direction, $y$ component of magnetisation $m_{y}$ remains zero always.As we decrease the temperature $m_{x}$ grows in each of the layers, resulting a thermal phase transition of the total system at a certain temperature $T_{c}$. This is shown in Figure-7. Due to the antiferromagnetic coupling between the adjacent layers, magnetic ordering are opposite in the respective layers. As a result, at a certain temperature ($T_{comp}$) they cancel each other, total magnetization becomes zero. But the individual layers possess a net magnetic moment. The compensation temperature $T_{comp}$ also increases (with small variation) as the strength of anisotropy increases.
@@
To visualize the spin ordering, we also study the spin morphology at different temperature for anisotropy $D=2.0$. This is depicted in Figure-8.At $T>>T_{c}$ the spins are randomly oriented at all layers, resulting in zero magnetisation. Near the critical temperature $T_{c}$ , thermal fluctuations are still strong enough to disrupt complete spin alignment, but a tendency emerges for spins to align along either the positive or negative $x$ direction. This marks the onset of magnetic domain growth, where regions of correlated spins begin to form and compete.
At the compensation point $T_{comp}$, the system exhibits a striking balance: the spins in the two outer layers align in one direction, while those in the middle layer orient in the opposite direction. As a result, the net magnetization of the entire system cancels out to zero, even though each layer retains a finite magnetization. This phenomenon is characteristic of ferrimagnetic-like ordering, where sublattice magnetizations compensate each other.
@@
At very low temperatures, the system reaches a fully ordered state. In this regime, the spins in the two outer layers are aligned along the negative $x$-direction, while the spins in the middle layer are aligned along the positive x-direction. This configuration reflects the combined effects of interlayer exchange interactions and anisotropy. To further elucidate the role of anisotropy in shaping the spin morphology, we also examined the angular distribution of spins for the case of $D=2.0$, where the easy-axis energetically favors spin orientations along $\theta=0$ or $\theta=\pi$. This is shown in Figure- . At high temperature, the majority of spins in all layers are distributed near $\theta=0$, and $\theta=\pi$ reflecting the dominant easy-axis alignment imposed by the positive anisotropy. Upon cooling, the distribution becomes progressively more ordered, with the relative populations $\theta=0$ and $\theta=\pi$ starting to compete. Near the compensation temperature, a distinct redistribution occurs: layers 1 and 3 exhibit a stronger spin population at $\theta=\pi$, whereas layer 2 retains it alignment at $\theta=0$. This antiparallel arrangement of sublattices lead to a vanishing total magnetisation while preserving finite sublattice magnetisation. At very low temperature, the system reaches a nearly complete alignment where layers 1 and 3 are oriented along $\theta=\pi$ and layer 2 aligned along $\theta=0$. At sufficiently low temperatures, the system recovers a finite total magnetisation as thermal fluctuation subside. Thus, while both the isotropic ($D=0$) and anisotropic ($D=2.0$) systems exhibit a compensation behaviour governed by the redistribution of opposing spin populations, the positive anisotropy merely selects preferred axes ($\theta=0,\pi$) without changing the underlying mechanism.
@@ 3.3 Finite size analysis
Finite-size effects play a crucial role in accurately characterising the thermodynamic phase transitions in numerical simulations. To investigate the finite-size behaviour of the clock model, we performed simulations for different system sizes $L=16,32,64,128,256$. The fourth-order Binder-cumulant $U_{L}(T)$, magnetisation $M(T)$ and corresponding susceptibilty $\chi(T)$ are studied as a function of temperature. The numerical results are shown in Figure-10 for the considered system sizes. Figure-10(a) illustrates the Binder Cumulant $U_{L}(T)$ where a well-defined intersection provides a robust indicator of the the critical temperature, as it is largely independent of $L$ near the thermodynamic transition. Figure-10(b) shows the magnetization, revealing the expected order-disorder behaviour as the temperature crosses the transition. The susceptibility curves for different system sizes highlights two distinct features: a prominant peak at the thermodyamic critical temperature and a secondary peak at the compensation point. From the Binder cumulant analysis, the intersection point of different $L$ curves yields a critical temperature $T_{c}=1.855$ marking the thermodynamic phase transition at anisotropy $D=2.0$. The susceptibility peak at $T_{c}$ exhibits a systematic increase with increasing system sizes, confirming the growth of critical correlation for the second-order phase transition. The susceptibility diverges in the thermodynamic limit ($L\rightarrow\infty$) according to the scaling $\chi_{max}\sim L^{\frac{\gamma}{\nu}}$. In contrast, the secondary peak associated with the compensation point does not show such scaling behaviour; its height remains nearly constant across the studied sizes. This suggest that the compensation phenomena is not a critical transition, but rather a finite, non-divergent reorganization of the magnetic sublattices.The finite size analysis confirms that the clock model undergoes a thermodynamic phase transition at $T_{c}$ , where critical fluctuations become system-size dependent, while the compensation point reflects a non-critical feature, a thermal anomaly.

### v2

@@ 3.2 Anisotropic regime: $D>0$
In this section, we explore the response of clock spins to the single site anisotropy. By varying the anisotropy parameter D, we systematically investigate the thermodynamics of the key observables such as magnetisation, susceptibility and specific heat. Figure-5(a) shows the thermal variation of total magnetisation as function of temperature for different anisotropy values. As temperature decreases, magnetisation grows, exhibiting a clear phase transition at $T_{c}$. Below $T_{c}$, the total magnetisation again vanishes at a lower temperature, denoted as the compensation point $T_{comp}$, indicating the compensation behaviour as already observed in the isotropic case. The critical temperature $T_{c}$ and compensation temperature $T_{comp}$ are precisely obtained from the peak positions of magnetic susceptibility curves Figure-5(b). The corresponding specific heat profiles further corroborate these findings by exhibiting peaks near both temperatures.
@@
Now, if we increase anisotropy D, the position of the maximum of the susceptibility shifts toward the right (higher temperature), indicating the critical temperature $T_{c}$ increases as $D$ increases. The compensation point $T_{comp}$ also shifts slightly. This is captured in the phase diagram in DTD-T plane shown in Figure-6. This diagram summarises the evolution of both the thermodynamic phase transition temperature $(T_{c})$ and the compensation temperature $(T_{comp})$ as a function of anisotropy $D$. A monotonic increase in $T_{c}$ is observed with increasing $D$, indicating that a stronger single-site anisotropy stabilizes the ordered phase and shifts the thermodynamic transition to higher temperatures. In contrast, $T_{comp}$ shows only a slight upward trend with increasing $D$ and remains well separated from $T_{c}$ across the probed range. The critical temperature $T_{c}$ is prescribed by collective critical fluctuations, which are highly sensitive to the anisotropy parameter $D$. Increasing $D$ enhances spin alignment and restrains thermal disorder, thereby stabilizes the ordered phase and shifts $T_{c}$ to higher temperatures. Besides, the compensation temperature $T_{comp}$ does not correspond to bonafide thermodynamic singularity. Instead, it arises from a subtle gross cancellation among the magnetisations of individual sublattices. This balance depends primarily on competing the intra and inter-layer exchange interactions and local spin configurations rather than on long-range critical fluctuations, resulting in only a weak variation of $T_{comp}$ with D.
@@
We also analyse the sublattice magnetisation of different layers for different strengths of anisotropy (Figure-7). As positive $D$ influences magnetic ordering along $x$ direction, $y$ component of magnetisation $m_{y}$ remains zero always. As we decrease the temperature, $m_{x}$ grows in each of the layers, resulting in a thermal phase transition of the total system at a certain temperature $T_{c}$. This is shown in Figure-7. Due to the antiferromagnetic coupling between the adjacent layers, magnetic orderings are opposite in the respective layers. At a certain temperature ($T_{comp}$), they cancel each other, leading to a net magnetisation of zero (even with nonzero sublattice magnetisation). The compensation temperature $T_{comp}$ also increases (with small variation) as the strength of anisotropy increases.
@@
The spin morphology at different temperatures for anisotropy $D=2.0$ is shown in Figure-8. At high temperature $T>>T_{c}$ the spins are randomly oriented at all layers, resulting in zero magnetisation. Due to the positive anisotropy majority of spins are aligned along 0 or $\pi$. Near the critical temperature $T_{c}$, thermal fluctuations are still strong enough to disrupt complete spin alignment, but a tendency emerges for spins to align along either the positive or negative $x$ direction. This marks the onset of magnetic domain growth, where regions of correlated spins begin to form and compete. At the compensation point $T_{comp}$, the system exhibits a striking balance: the spins in the two outer layers align in one direction, while those in the middle layer orient in the opposite direction. As a result, the net magnetisation of the entire system cancels out to zero, even though each layer retains a finite magnetisation. This phenomenon is characteristic of ferrimagnetic-like ordering, where sublattice magnetisations compensate each other. At very low temperatures, the system reaches a fully ordered state. In this regime, the spins in the two outer layers are aligned along the negative $x$-direction, while the spins in the middle layer are aligned along the positive $x$-direction. This configuration reflects the combined effects of interlayer exchange interactions and anisotropy.
@@
To further elucidate the role of anisotropy in shaping the spin morphology, we also examined the angular distribution of spins for the case of $D=2.0$, where the easy-axis energetically favours spin orientations along $\theta=0$ or $\theta=\pi$. This is shown in Figure-9. At high temperature, the majority of spins in all layers are distributed near $\theta=0$, and $\theta=\pi$ reflecting the dominant easy-axis alignment imposed by the positive anisotropy. Upon cooling, the distribution becomes progressively more ordered, with the relative populations $\theta=0$ and $\theta=\pi$ starting to compete. Near the compensation temperature, a distinct redistribution occurs: layers 1 and 3 exhibit a stronger spin population at $\theta=\pi$, whereas layer 2 retains it alignment at $\theta=0$. This antiparallel arrangement of sublattices leads to a vanishing total magnetisation while preserving finite sublattice magnetisation. At very low temperature, the system reaches a nearly complete alignment where layers 1 and 3 are oriented along $\theta=\pi$ and layer 2 aligned along $\theta=0$. At sufficiently low temperatures, the system recovers a finite total magnetisation as thermal fluctuations subside. Thus, while both the isotropic ($D=0$) and anisotropic ($D=2.0$) systems exhibit a compensation behaviour governed by the redistribution of opposing spin populations, the positive anisotropy merely selects preferred axes ($\theta=0,\pi$) without changing the underlying mechanism.
@@ 3.3 Finite size analysis
Finite-size effects play a crucial role in accurately characterising the thermodynamic phase transitions in numerical simulations. To investigate the finite-size behaviour of the clock model in anisotropic regime ($D=2.0$), we performed simulations for different system sizes $L=32,64,128,256$. The fourth-order Binder-cumulant $U_{L}(T)$, magnetisation $M(T)$ and corresponding susceptibility $\chi(T)$ are studied as a function of temperature. The numerical results are shown in Figure-10 for the considered system sizes. Figure-10(a) illustrates the Binder Cumulant $U_{L}(T)$ where a well-defined intersection provides a robust indicator of the critical temperature, as it is largely independent of $L$ near the thermodynamic transition. Figure-10(b) shows the magnetisation, revealing the expected order-disorder behaviour as the temperature crosses the transition. The susceptibility curves for different system sizes manifest two salient features: a prominent peak at the thermodynamic critical temperature and a secondary peak at the compensation point. From the Binder cumulant analysis, the intersection point of different $L$ curves yields a critical temperature $T_{c}=1.855$ marking the thermodynamic phase transition at anisotropy $D=2.0$. The susceptibility peak at $T_{c}$ exhibits a systematic increase with increasing system sizes, confirming the growth of critical correlation for the second-order phase transition. The susceptibility diverges in the thermodynamic limit ($L\rightarrow\infty$) according to the scaling $\chi_{max}\sim L^{\frac{\gamma}{\nu}}$. In Figure-11, we present maxima of susceptibility as function of system sizes in logarithmic scale. The straight line fit estimates the critical exponent ratio $\frac{\gamma}{\nu}=0.532$. In contrast, the secondary peak associated with the compensation point does not show such scaling behaviour; its height remains nearly constant across the studied system sizes. This suggests that the compensation phenomena is not a critical transition, but rather a finite, non-divergent reorganisation of the magnetic sublattices. The finite-size analysis confirms that the clock model undergoes a thermodynamic phase transition at $T_{c}$ , where critical fluctuations become system-size dependent, while the compensation point reflects a non-critical feature, inherent to the system’s response.